OpenStudy (anonymous):

A 2-digit number is divisible by both 2 and 5. Half this number is less than 6 times the tens digit of the number by 4. Find this 2-digit number. @amistre64

OpenStudy (amistre64):

the LCM of 2 and 5 is 10, i wonder if its a multiple of 10

OpenStudy (amistre64):

"6 times the tens digit of the number by 4" cant make sense of that part

OpenStudy (anonymous):

half the number is 4 less than 6 times the tens digit of the number?

OpenStudy (anonymous):

i cannot get the idea haha

OpenStudy (anonymous):

i just type this from the paper

OpenStudy (anonymous):

grind it till you find it you only have 9 choices

OpenStudy (amistre64):

Half this number is less than 6 times the tens digit of the number by 4 n/2 + 4 = 6t maybe?

OpenStudy (anonymous):

the two-digit number is 10*a + b b is zero because only 0 is a single digit and divisible by 2 and 5 5a+4 = 6a

OpenStudy (anonymous):

anyone see what I did there?

OpenStudy (anonymous):

@amistre64 i don't know but i think so

OpenStudy (anonymous):

@Peter14 i got it

OpenStudy (anonymous):

lets try 20: the tens digit is 2 half of it is 10 \(10=6\times 2-4=12-4=8\) nope, not \(20\)

OpenStudy (amistre64):

10t/2 + 4 = 6t 5t + 4 = 6t 4 = t maybe its 40?

OpenStudy (anonymous):

@Peter14 but why 5a?

OpenStudy (anonymous):

the answer is 40

OpenStudy (anonymous):

because it's half of 10a

OpenStudy (anonymous):

i don't know the stpes...

OpenStudy (anonymous):

@Peter14 okay..

OpenStudy (anonymous):

experiment! not every thing needs to be solved via an equation check and see how it works, then you will know the answer

OpenStudy (anonymous):

"Half this number is less than 6 times the tens digit of the number by 4" reword this as: half this number is 4 less than 6 times the ten digit of the number

OpenStudy (anonymous):

oh i see. thank you @satellite73 @Peter14 @amistre64 i am still confusing the sentence that @Peter14 has just mentioned

OpenStudy (amistre64):

its a confusing sentence, i agree. but i finally read it as Peter defined it

OpenStudy (anonymous):

the sentence boils down to this: (10*a + b)/2 = 6*a - 4 since we already know b=0, this simplifies to: 5*a = 6*a-4

OpenStudy (anonymous):

oh i really got it . thank you three very much

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

i am thinking who should i give the medal to ..